The population P, t years after the initial observation is given by the formula:

pt=(100000)/(2+3e-0.05t)

Determine the exact size of the population and the time in years when the rate of growth is maximum


1
Expert's answer
2022-02-22T08:35:25-0500
"P(t)=\dfrac{100000}{2+3e^{-0.05t}}"

"P'(t)=\dfrac{15000e^{-0.05t}}{(2+3e^{-0.05t})^2}=\dfrac{15000e^{0.05t}}{(3+2e^{0.05t})^2}"

"P''(t)=\dfrac{750e^{0.05t}(3-2e^{0.05t})}{(3+2e^{0.05t})^3}"

"P''(t)=0=>\dfrac{750e^{0.05t}(3-2e^{0.05t})}{(3+2e^{0.05t})^3}=0"

"e^{0.05t}=1.5"

"t=20\ln 1.5"

The rate of growth is maximum at "t=20\ln 1.5."

Then


"P(20\ln 1.5)=\dfrac{100000}{2+3(2/3)}=25000"

The exact size of the population is 25000 and the time is "20\ln 1.5\approx8.11" years when the rate of growth is maximum.